OpenStudy (anonymous):

What is the soulution of (9^x)-(3^x+1)-28=0

OpenStudy (anonymous):

3^2x - 3^x -29=0

OpenStudy (anonymous):

Thanks again

OpenStudy (anonymous):

actually 6x-29=0 which means x=29/6

OpenStudy (anonymous):

oh shooo forget it